OpenStudy (anonymous):

How wold I solvle the following system of equations. 2x + y = 3 x = 2y - 1 ?

OpenStudy (anonymous):

okay so get the y alone on the first equation so it would be y=-2x+3 -2y = -x-1 then multiply the first equation by 2 so you can cancel both y's 2y=-4x+6 -2y=-x-1 which when you add them together becomes 0=-5x+5 when you subtract 5x and solve, x=1

OpenStudy (anonymous):

the last time i did Maths was back in 2010 but i believe a Q of this nature can be easily solved simultaneously. 2x + y= 3 you must shift your eqeation by making one variable a subject of your formula. if we say, y= -2x + 3 then continue to solve both equations @ the very same time by repclacing where there is y by your 2nd equation. now we can then use 2x + (-2x + 3) = 0, then 2x -2x + 3 = 0 and by adding you obtain 3= 0 coz 2x-2x canceles out. but if we take 2x to be our subject of formula we can go about and do as, 2x + y = 3 2x= 3 - y x= -y/2 +3/2 now we can say from our original equation 2(-y/2 + 3/2) +y = 3 -y + 3 + y = 3

OpenStudy (radar):

The problem did not say what method to use to solve this system. But the equations provided would make it very simple to use the substitution method. The second equation x = 2y - 1 already has the x isolated so just substitute this for x in the first equation: 2(2y - 1) + y = 3 Note how the 2y - 1 from the 2nd equation is substituted for x in 1st. Now solve for y 4y - 2 + y = 3 5y = 5 y = 1 Now substitute the value "1" for y in either one of the original equations and solve for x.

OpenStudy (anonymous):

Thank you everyone for the help. I am looking to find the graphing coordinates.

OpenStudy (anonymous):

geometric figure (point) x-1